OpenStudy (amistre64):

Find the volume of the solid created by the bounded region of (x=y-y^2) and (x=y^2 -3) when it is spun around the line (x=-4).

OpenStudy (anonymous):

Ah this question still o.O

OpenStudy (amistre64):

Yeah :) but I found out what I was doing wrong.... other than trying to drink and derive :)

OpenStudy (anonymous):

So we at least know you can add 4 to all terms to change the bounds and have it revolved around x = 0

OpenStudy (anonymous):

And it's a torus, right? So you need to integrate the area of the torus around the line in a full circle.

OpenStudy (anonymous):

A cross section of the torus.

OpenStudy (amistre64):

yep...

OpenStudy (anonymous):

Can't you just integrate it?

OpenStudy (amistre64):

its pi (S) [4+F(y)]^2 - [4+G(y)]^2 | -1, 3/2 I kept trying: pi (S) [(4+F(y)) - (4+G(y))]^2 | -1, 3/2

OpenStudy (amistre64):

math: yes, but couldnt figure out how to last night :)

OpenStudy (anonymous):

\[x = y - y^2 + 4\] \[x = y^2 + 1\]

OpenStudy (anonymous):

Okay so the bounds are y = -3, y = 1/2

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

What?

OpenStudy (amistre64):

Krb: goog good. the bounds of integration are -1 and 3/2

OpenStudy (anonymous):

Yeah those are the bounds ^

OpenStudy (anonymous):

i need to figure out how to use the equations to find the solids of revolution..if this is x=-4, then the line we get is a vertical one right? so its been revolved upwards?

OpenStudy (amistre64):

you got the equation right; now square them individually before you subtract them..

OpenStudy (anonymous):

So \[\int\limits_{(-1)}^{(3/2)}(y - y^2 + 4) - (y^2 + 1)\]

OpenStudy (anonymous):

That's what i would think

OpenStudy (anonymous):

Krb you're always on this site. I like how when I look at paul's notes, you were the first one on open study.

OpenStudy (amistre64):

think; the rotation is around the "y axis" so to speak

OpenStudy (anonymous):

yeah..thats what i mean, so its like integrating the difference of squares of the equations right? and these equations signify -----?

OpenStudy (amistre64):

Krb: getting closer. we need to integrate [F(y)]^2 - [G(x)]^2 in the bounds

OpenStudy (anonymous):

Yes, normally when integrating with respect to x you are finding the are underneath the curve that drops to the x axis. So in this case you are finding the area "to the side of the curve" that drops to the y axis

OpenStudy (amistre64):

Think; exactly :)

OpenStudy (anonymous):

I'm confused as to why it would be squared

OpenStudy (amistre64):

typo...[G(y)]^2

OpenStudy (anonymous):

i guess it should be squared...shouldn't it?

OpenStudy (amistre64):

Krb: if we do it seperately for each on, there is no real "area" so to speak until we combine them. Does that make sense?

OpenStudy (amistre64):

pi* (S) F(y)^2 dy - (S) G(y)^2 dy The equations ARE our Radiusessess

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

if y1 & y2 are the curves then , \[\int\limits_{}^{}[y1^{2}-y2^{2}]\] gives the volume of the solid? where y1=upper curve & y2 the lower

OpenStudy (anonymous):

multiplied by a pi

OpenStudy (anonymous):

SO we are looking at it from a top down view, and then you need pi * r^2 for both "circles"

OpenStudy (amistre64):

exactly

OpenStudy (anonymous):

on efor inner and the other for outer 'circle' i guess

OpenStudy (amistre64):

think: thats the gyst of it yeah

OpenStudy (anonymous):

thnx!!

OpenStudy (anonymous):

\[[\pi \int\limits_{-1}^{(3/2)}(y - y^2 + 4)^2] - [\pi \int\limits_{-1}^{(3/2)}(y^2 + 1)^2]\]

OpenStudy (amistre64):

Krb: your a genius :)

OpenStudy (anonymous):

No, you are :)

OpenStudy (anonymous):

u both are :)

OpenStudy (amistre64):

we should start a mensa club right here lol

OpenStudy (anonymous):

the open study math mensa

OpenStudy (anonymous):

:) :) so krb ur'e pursuing your masters too?

OpenStudy (amistre64):

C1 - C2 is what we were missing last night. 2pi(R) - 2pi(r) does not equal 2pi(R-r)

OpenStudy (amistre64):

it does, but it doesnt when you integrate lol

OpenStudy (anonymous):

wer those constants that u'll missed out yesterday?

OpenStudy (anonymous):

Ends up being 85.9029

OpenStudy (anonymous):

think: I'm honestly not sure

OpenStudy (anonymous):

krb: that's alright

OpenStudy (anonymous):

i dunno why it gets posted twice always, when i click only once!

OpenStudy (anonymous):

I think I probably would like to get a Masters

OpenStudy (anonymous):

good